OpenStudy (anonymous):

Please help! Solve for X 2/5(12+x)=6-1/4(9-x) I'm not sure what to do first...

OpenStudy (shadowlegendx):

\[\frac{ 2 }{ 5 }(12 + x) = 6 - \frac{ 1 }{ 4 }(9-x)\]

OpenStudy (anonymous):

yes, sorry I don't know how to properly express the problem using my keyboard

OpenStudy (shadowlegendx):

First lets distribute \[\frac{ 2 }{ 5 }(12 + x) \]

OpenStudy (shadowlegendx):

I used the equation button located here: http://prntscr.com/4kzu9k

OpenStudy (anonymous):

ok thank you I will use it in the future. so distributing would it be like \[\frac{ 2 }{ 5 }\times \frac{ 12 }{ 1 }\] first?

OpenStudy (shadowlegendx):

\[4.8 + \frac{ 2 }{ 5 }x = 6 - 2.25 - \frac{ 1 }{ 4 }x\] This is what you should have gotten after successfully distributing on both sides

OpenStudy (shadowlegendx):

 This is distributing. Multiplying the 1 by both the 2 and 3 inside The result would be 2 + 3